# Fast chaos indicator from auto-differentiation for dynamic aperture optimization

### v1

@@ 1 INTRODUCTION
Other efficient methods, such as frequency map analysis (FMA) and the reversibility error method (REM), have also been studied in the accelerator community to accelerate DA estimation[10, 11, 12, 13, 14, 15, 16, 17]. In the FMA approach, changes in the particle trajectory frequency, quantified as the tune diffusion rate, are employed as indicators of chaotic motion. A large tune diffusion rate corresponds to strongly chaotic behavior. In the REM approach, a particle is tracked forward for a specified number of turns and then tracked backward to its initial time or longitudinal coordinate. The separation between the initial and final phase‑space coordinates provides a measure of sensitivity to perturbations from numerical round‑off errors. For regular trajectories, this separation grows as a power law with respect to the number of steps, whereas for chaotic trajectories it grows exponentially.
@@ 2 Forward mode automatic differentiation
The automatic differentiation (AD) module used in this study is based on the Truncated Power Series Algebra (TPSA) method but is limited to first derivatives [26]. This restriction substantially reduces the computational complexity of the arithmetic operations and decreases execution time by orders of magnitude, while also simplifying the module’s implementation. In the artificial intelligence and machine learning (AI/ML) community, this approach is equivalent to the dual‑number implementation of forward‑mode AD [27].
@@ 4 Test of the Indicator Using a Hénon–Heiles Potential
where the Lie operator :H::H: is defined by :H:g=[H,g]:H:g=[H,g] for an arbitrary function $g$.
@@ 5 Application to Dynamic Aperture Optimization
In this study, we use one version of the ALS-U lattice design for demonstration purposes; it does not represent the latest official ALS-U configuration. Particle tracking is performed using the differentiable tracking code JuTrack [25] to determine the dynamic aperture. Figure 2 compares the horizontal, vertical beta functions, and horizontal dispersion evolution through one of $12$ cells of the nominal ALS-U lattice obtained using JuTrack and the Elegant [31] simulation code. The results show excellent agreement between the two codes. The nominal working tunes for the design are $41.374$ (horizontal) and $20.379$ (vertical).
@@
Along the $x$-axis, the $y$ coordinate is scanned from zero until the log Frobenius norm exceeds a threshold value of one. The area beneath the resulting boundaries defines the estimated dynamic aperture. To encourage symmetry, areas for $x>0$ and $x<0$ are computed separately, and the absolute difference between the two is subtracted from the total area $A$ as a penalty. The objective function is therefore
@@
To validate the optimization results, we performed direct particle tracking for $1000$ turns using the optimized quadrupole settings to obtain the dynamic aperture boundary. Figure 6 compares the boundaries obtained from the nominal and optimized lattices. The optimized configuration yields a noticeably larger dynamic aperture, consistent with the predictions of the tangent map norm indicator.

### v2

@@ 1 INTRODUCTION
Other efficient methods, such as frequency map analysis (FMA) and the reversibility error method (REM), have also been studied in the accelerator community to accelerate DA estimation[10, 11, 12, 13, 14, 15, 16, 17]. In the FMA approach, changes in the particle trajectory frequency, quantified as the tune diffusion rate, are employed as indicators of chaotic motion in the presence of high-order multipole nonlinear fields. A large tune diffusion rate corresponds to strongly chaotic behavior; in such cases, particles within chaotic regions are typically lost quickly in a lattice with misalignment field errors. In the REM approach, a particle is tracked forward for a specified number of turns and then tracked backward to its initial time or longitudinal coordinate. The separation between the initial and final phase‑space coordinates provides a measure of sensitivity to perturbations from numerical round‑off errors. For regular trajectories, this separation grows as a power law with respect to the number of steps, whereas for chaotic trajectories it grows exponentially.
@@ 2 Forward mode automatic differentiation
The automatic differentiation (AD) module used in this study is based on the Truncated Power Series Algebra (TPSA) method [26] but is limited to first derivatives [27]. This restriction substantially reduces the computational complexity of the arithmetic operations and decreases execution time by orders of magnitude, while also simplifying the module’s implementation. In the artificial intelligence and machine learning (AI/ML) community, this approach is equivalent to the dual‑number implementation of forward‑mode AD [28].
@@ 4 Test of the Indicator Using a Hénon–Heiles Potential
where the Lie operator :H::H: is defined by :H:g=[H,g]:H:g=[H,g] for an arbitrary function $g$ [31].
@@ 5 Application to Dynamic Aperture Optimization
In this study, we use one version of the ALS-U lattice design for demonstration purposes; it does not represent the latest official ALS-U configuration. Particle tracking is performed using the differentiable tracking code JuTrack [25] to determine the dynamic aperture. JuTrack is a modern, AD-enabled accelerator tracking software package written in the Julia programming language. It has been thoroughly benchmarked against widely used accelerator codes, such as MADX, demonstrating consistent accuracy in particle tracking and optics functions. Furthermore, JuTrack provides accurate derivatives with respect to any pre-defined simulation parameters through automatic differentiation. These AD results have been fully validated against finite difference methods. Figure 2 compares the horizontal, vertical beta functions, and horizontal dispersion evolution through one of $12$ cells of the nominal ALS-U lattice obtained using the first-order TPSA method in the JuTrack and the Elegant [33] simulation code. The results show excellent agreement between the two codes. The nominal working tunes for the design are $41.374$ (horizontal) and $20.379$ (vertical). In addition to the optics function benchmarks, we compared single-particle tracking results from both codes. Figure 3 shows the element-by-element evolution of the transverse phase space through the accelerator using JuTrack and Elegant. As demonstrated, both codes show excellent agreement.
@@
Figure 3: Element by element evolution of the $X$ (top left), $X^{\prime}$ (top right), $Y$ (bottom left) and $Y^{\prime}$ (bottom right) as computed with JuTrack and Elegant for a particle starting at coordinates $(0.001,0,0.001,0,0,0)$.
@@
Along the $x$-axis, the $y$ coordinate is scanned from zero until the log Frobenius norm exceeds a threshold value of one. The choice of a threshold value of one is based on the rapid color change shown in Figure 4 and is guided by the shape of the dynamic aperture from the 1,000-turn direct tracking. We also performed an optimization using a threshold value of 1.2 instead of one. The resulting optimized parameters were not very sensitive to this change in the threshold value. The area beneath the resulting boundaries defines the estimated dynamic aperture. To encourage symmetry, areas for $x>0$ and $x<0$ are computed separately, and the absolute difference between the two is subtracted from the total area $A$ as a penalty. The objective function is therefore
@@
In the above optimization, all sextupoles in the accelerator were kept constant. To explore further improvements, we included two sextupole families located outside the dispersion region alongside the two quadrupole families. The ranges for these two sextupole families were set to [1, 7] $m^{-3}$ and [-1020, -840] $m^{-3}$, respectively. The resulting optimized dynamic aperture area is $1.42\times 10^{-6}m^{2}$, achieved with the optimal settings of (13.84,13.77,4.75,930.92)(13.84,-13.77,4.75,-930.92). Ultimately, including these two sextupole families in the optimization did not significantly improve the dynamic aperture area.
To validate the optimization results, we performed direct particle tracking for $1000$ turns using the optimized quadrupole settings (optimal solution 1) and quadrupole and sextupole settings (optimal solution 2) to obtain the dynamic aperture boundary. Figure 7 compares the boundaries obtained from the nominal and optimized lattices. The optimized configuration yields a noticeably larger dynamic aperture, consistent with the predictions of the tangent map norm indicator.
@@ 6 Conclusions
The norm of the one-turn tangent map obtained from automatic differentiation can serve as an effective indicator for fast initial optimization. However, these initial optimized parameters should be verified or further refined using more accurate, albeit slower, methods such as long-term direct or differentiable tracking, or MFA.
Furthermore, other effective methods, such as analytical-based resonance correction [37, 38] and direct tracking with a flood-fill pattern [39], are also used for dynamic aperture optimization. The method proposed in this paper can be viewed as an alternative approach to these techniques.